Identical charges separated by 1 cm experience a repulsion of 0.1 N. Find the charges.
Answer #1

Force = kq2/r2

r = 1 cm = 0.01 m

k = 8.9*109 N m2 / C-2

0.1 = 8.9*109 *q2 / (0.01)2

=> q2 = 0.1 * (0.01)*(0.01)/ 8.9*109

Charge => q =  3.3 * 10-3 C
Answer #2

F = kQq / d
Answer #3


given q1 = q2 = q=?

d = 0.01m....
F = 0.1N...
F = k*q1*q2/d^2...
F = k*q^2/d^2...

q^2 = F*d^2/k = 0.1*0.01^2/(9*10^9) = 0.112*10^-14

q = 0.334*10^-7 C = 0.0334 uC = 33.4 nC